Q1.A 50 mm diameter stream of water strikes a 0.5 m by 0.5 m square plate which is at an angle of 30 degrees with the stream’s direction. The velocity of the water in the stream is 20 m/s and the jet strikes the door at its centre of gravity. If the plate is hinged at A, by neglecting the friction and self-weight of the plate, determine the normal force P applied at the upper edge of the plate to keep the plate in equilibrium. You may assume that the velocities of the jet remain unchanged after split.
1
Expert's answer
2019-04-08T09:17:24-0400

The normal force P:


"P=F \cos30° (1)"

Newton’s second law of motion:


"\u0394mv=F\u0394t (2)"

where Δm is the change of water mass


The change of water mass:


"Δm=ΔV\times \rho =SvΔt\rho =\frac {\pi d^2}{4} vΔt\rho (3)"



Put (3) in (2):


"\frac {\pi d^ 2}{4} vΔt\rho\times v= FΔt (4)"

We got:


"F=\frac {\pi d^ 2}{4} v^2 \rho (5)"

where ρ is the water density

Put (5) in (1):


"P=\frac {\pi d^ 2}{4} v^2 \rho \cos30° (6)"

Using (6) we got:

P=68.3 N


Answer:

68.3 N